OpenStudy (adamaero):

To provide for a college education for her son, a woman opened an account in which equal annual deposits were made. The first deposit was made on January 1, 1998, and the last deposit was made on January 1, 2015. The yearly college expenses including tuition were estimated to be $9000, for each of the 4 years. Assuming the interest rate to be 4.5%, how much did the mother have to deposit each year in the account for the son to draw $9000 per year for 4 years beginning January 1, 2015. (Answer: $1255)

OpenStudy (adamaero):

I did Deposits = Withdrawals A(P/A, 4.5, 18) = 9000(P/A, 4.5, 4)(P/F, 4.5, 18) then solved for A...but I got $1202--where did I go wrong?!

OpenStudy (anonymous):

i have no idea what formula does this do you?

OpenStudy (adamaero):

It's two compound interests formulas, taken from a table: http://global.oup.com/us/companion.websites/9780199778126/pdf/Appendix_C_CITables.pdf#page=14

OpenStudy (adamaero):

A(12.6) = 9000(3.588)(.4528)